# Variable Elliptic Structures on the Plane: Transport Dynamics, Rigidity, and Function Theory

### v2

@@ Conclusion
This monograph developed a variable function theory on planar domains in which the imaginary unit is no longer a fixed algebraic constant but a geometric object that varies with position. The guiding principle throughout has been that the symbol $i$ should be treated as intrinsic data: a moving generator of a rank–two real algebra bundle whose defining structure reduction
$i^{2}+\beta(x,y)\,i+\alpha(x,y)=0$
is allowed to vary. Once this perspective is adopted, the derivatives $i_{x}$ and $i_{y}$ are no longer formal artifacts but genuine geometric quantities, and their algebraic resolution becomes the natural starting point of the theory.
A central object emerging from this viewpoint is the intrinsic combination
$i_{x}+i\,i_{y}.$
This quantity appears prior to any analytic notion of holomorphicity. It governs the variability of the structure itself and, as a consequence, controls both the inhomogeneity of the generalized Cauchy–Riemann system and the failure of the Leibniz rule for the operator $\partial_{\bar{z}}$. These phenomena are not independent: they are different manifestations of the same structural obstruction produced by differentiating the moving generator.
A fundamental outcome of the analysis is that the evolution of the structure coefficients $(\alpha,\beta)$ is governed by a first–order transport law that is already present at the structural level. By differentiating the structure reduction and eliminating the derivatives of $i$, the coefficient dynamics admits a reformulation in terms of the spectral parameter
@@
leading, in full generality, to a (possibly forced) complex Burgers equation written in the standard complex algebra. This transport formulation is intrinsic and unavoidable: it reflects the fact that allowing the elliptic structure to vary forces its coefficients to evolve along characteristic curves.
Rigidity occupies a precise and limited position within this framework. It is not the source of the Burgers dynamics, but rather the distinguished conservative regime of the universal transport law. The vanishing condition
@@
selects exactly those variable elliptic structures for which the Burgers equation is unforced. In this conservative case, several obstructions collapse simultaneously: the generalized Cauchy–Riemann operator becomes a derivation, the real Cauchy–Riemann system becomes homogeneous, and the generator $i$ is transported compatibly with its own multiplication law. Rigidity is therefore not an auxiliary hypothesis but a compatibility condition that restores integrability within an already dynamical setting.
@@
The explicit $\varepsilon$–family of rigid elliptic structures illustrates this conservative regime in concrete terms. It shows that rigidity does not force triviality: genuinely nonconstant structures exist, though ellipticity naturally restricts their domain of definition. From the transport viewpoint, these examples reflect the familiar behavior of first–order hyperbolic systems, where characteristic crossing limits global regularity and constrains smooth extensions.
On the analytic side, rigidity is exactly the condition under which an integral representation theory can persist. In the rigid regime, a Cauchy–Pompeiu formula adapted to variable elliptic structures is recovered. The natural Cauchy 11–form
$d\tilde{z}=dy-i(z)\,dx$
is not closed, and its non–closure produces an unavoidable interior correction term. Rather than being an artifact, this correction singles out a canonical first–order operator
$Df=\partial_{\bar{z}}f+\tfrac{1}{2}f\,i_{y},$
whose null space provides the appropriate notion of holomorphicity in the variable setting. Although covariantly holomorphic sections are not closed under the pointwise product, a scalar gauge transformation restores closure through a natural weighted multiplication, showing how analytic structure can be recovered by compensating for variability rather than suppressing it.
The scope of the present work is deliberately local and explicit. Even in the conservative regime, global questions remain subtle: the choice of a continuous generator $i(z)$ is typically local, gauge trivializations may exhibit monodromy, and the transport nature of the coefficient dynamics explains why globally smooth structures are highly constrained. Beyond rigidity, the analytic function theory necessarily acquires source terms, while the transport–controlled evolution of the coefficients remains intact.
Closing perspective. The central lesson of this work is that once the complex structure is allowed to move, its differentiation becomes a primary geometric invariant. Transport comes first, integrability second, and analytic closure emerges only as a distinguished limiting case. Rigid elliptic structures do not generate the dynamics; they represent the rare situation in which the universal transport law becomes conservative and the analytic and algebraic layers of the theory align. In this way, variable elliptic structures reveal a natural hierarchy in which geometry dictates dynamics, and dynamics, in turn, governs analysis.
Acknowledgements

### v3

@@ Conclusion
This monograph set out from a simple but decisive shift in viewpoint: the imaginary unit should not be treated as a fixed algebraic constant, but as intrinsic geometric data. Allowing the generator $i(x,y)$ of a rank–two real algebra bundle to vary forces the algebra itself to become part of the geometry of the plane. Once this step is taken, differentiation of the structure is no longer optional, and the derivatives $i_{x}$ and $i_{y}$ emerge as genuine geometric quantities rather than formal artifacts.
The central object organizing the theory is the intrinsic obstruction
$G:=i_{x}+i\,i_{y}.$
This quantity appears prior to any analytic notion of holomorphicity. It measures the incompatibility between differentiation in the base variables and multiplication in the fibers, and it governs, simultaneously and unavoidably, all deviations from the classical theory: the inhomogeneity of generalized Cauchy–Riemann systems, the failure of the Leibniz rule, and the appearance of source terms in first–order structure equations. These phenomena are not imposed by analytic assumptions; they are forced by the geometry of a moving elliptic algebra.
A first structural outcome of the theory is that variability of the coefficients $(\alpha,\beta)$ is not arbitrary. Eliminating the derivatives of the generator from the structure reduction leads to a universal transport law for the spectral parameter
@@
which takes the form of a (possibly forced) complex inviscid Burgers equation written in the standard complex algebra. This transport equation belongs to the structural layer of the theory: it arises solely from allowing the elliptic structure to vary and does not rely on integrability, coordinate trivialization, or function–theoretic input.
Within this universal transport framework, rigidity occupies a precise and limited role. The condition
@@
does not generate the dynamics; it selects the conservative regime of the already–existing transport law. In this regime, the Burgers equation becomes unforced, and several independent obstructions collapse simultaneously: the generalized Cauchy–Riemann operator becomes a derivation, the real first–order system becomes homogeneous, and the generator is transported compatibly with its own multiplication law. Rigidity is therefore a compatibility condition, not a simplifying assumption.
@@
The explicit $\varepsilon$–family constructed in this work demonstrates that rigidity does not force triviality. Genuinely nonconstant elliptic structures exist, although ellipticity naturally restricts the domain of definition. From the transport viewpoint, these restrictions reflect characteristic phenomena familiar from first–order hyperbolic systems and are not pathologies of the theory.
On the analytic side, rigidity is exactly the threshold at which a coherent function theory re–emerges. In the rigid regime, a Cauchy–Pompeiu formula adapted to variable elliptic structures persists. The natural Cauchy 11–form is not closed, and its non–closure produces an unavoidable interior correction term. Rather than being an artifact, this correction identifies a canonical covariant Cauchy–Riemann operator whose null space provides the appropriate notion of holomorphicity in the variable setting. Although covariantly holomorphic sections are not closed under pointwise multiplication, a scalar gauge transformation restores closure through a natural weighted product. Power series representations follow from this framework, but only after freezing the algebra at the expansion point, making explicit where classical arguments genuinely fail in the variable case.
Finally, the theory interfaces naturally with the classical Vekua program. General first–order elliptic systems induce variable elliptic structures, and their reduction to Vekua form reveals the intrinsic obstruction as the true geometric source of inhomogeneity. Rigidization, formulated as an ordinary differential equation along characteristic curves, provides a constructive mechanism to pass from non–rigid to rigid structures without solving global Beltrami equations. In the rigid regime, sharp similarity principles, initial value problems, and algorithmic solution schemes become available, even in the presence of non–uniform ellipticity.
The scope of this work has been deliberately local and explicit. Rather than appealing to abstract integrability frameworks, the theory has been developed at the level of concrete first–order computations in order to keep the geometric and dynamical mechanisms fully visible. From this perspective, variable elliptic structures exhibit a clear hierarchy: transport is fundamental, rigidity is selective, and analytic closure is a distinguished consequence rather than an axiom.
This hierarchy reframes classical complex analysis as a special case of a broader geometric theory.
A fixed complex structure gives analysis. A variable one gives transport. Rigidity is where the two agree to dance.
Here, the algebra begins to move.